# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v1

@@ B.2 Proof of Lemma 11
$\overset{(b)}{\leq}\frac{1}{2}\exp\left(-\frac{C_{6}}{4}k\log k\right).$ (50)
Here, we apply the union bound in inequality (a), and invoke the fact that $C_{6}$ is large enough (e.g., $C_{6}>8C_{\mathsf{cover}}^{\prime}$) in equality (b).
Regarding the set $\mathcal{G}$, we analyze it in a similar manner. First, we recall some standard concentration bounds below; the proof can be found in Wainwright, (2019, Chapter 2).
@@
$\mathbb{P}(\|Z_{1}\|_{2}\geq t)\leq e^{-t^{2}/2}\qquad\text{and}\qquad\mathbb{P}(\|Z\|_{2}^{2}\geq\sqrt{d}+t)\leq e^{-t^{2}/2},$
@@
$\mathbb{P}(Z\notin\mathcal{G})$ $\overset{(a)}{\leq}\mathbb{P}\left(\|Z\|_{2}>\sqrt{d}+\sqrt{C_{6}k\log k}\right)+\sum_{i=1}^{s}\sum_{j=1}^{s}\mathbb{P}\left(\langle x_{i}^{*}-x_{j}^{*},Z\rangle>\sqrt{C_{6}k\log k}\|x_{i}^{*}-x_{j}^{*}\|_{2}\right)$
@@
$\overset{(b)}{\leq}\frac{1}{2}\exp\left(-\frac{C_{6}}{4}k\log k\right),$ (51)
@@
Since $X_{t}\overset{\mathrm{d}}{=}\sqrt{1-\sigma_{t}^{2}}X_{0}+\sigma_{t}Z$, combining inequalities (50) and (51) yields
@@
B.3 Proof of Lemma 12
For ease of our presentation, we fine it convenient to define the following set
$\mathcal{E}_{t}(x)\coloneqq\left\{x_{0}\,\Big{|}\,\big{\|}x_{0}-x_{i(x)}^{*}\big{\|}_{2}\geq\frac{\sigma_{t}}{\sqrt{1-\sigma_{t}^{2}}}\cdot\sqrt{Ck\log k}\right\},$
where $C>C_{7}$ for some sufficiently large constant $C_{7}>0$ (to be specified momentarily). It suffices to bound $\mathbb{P}(\mathcal{E}_{t}(x)\mid X_{t}=x)$. Denote by $q_{X_{t}|X_{0}}(\cdot|x_{0})$ the conditional density function of $X_{t}$ given $X_{0}=x$, and $q_{0}$ the measure induced by $X_{0}$. Making use of the Bayes formula, we can obtain
@@ B.3 Proof of Lemma 12
where the last inequality holds by exploiting $i(x)\in\mathcal{I}$ and the definition of $\mathcal{I}$ in (46a). The proof then boils down to upper bounding the exponential term above.
Given that $x\in\mathcal{T}_{t}$, we can decompose it as $x=\sqrt{1-\sigma_{t}^{2}}x_{0}^{\prime}+\sigma_{t}z$ for some $x_{0}^{\prime}\in\mathcal{B}_{i(x)}$ and $z\in\mathcal{G}$. For any $x_{0}\in\mathcal{E}_{t}(x)$ and $\widetilde{x}_{0}\in\mathcal{B}_{i(x)}$, suppose that $x_{0}\in\mathcal{B}_{j}$, then direct expansion gives
@@
$=-(1-\sigma_{t}^{2})\|x_{0}^{\prime}-x_{0}\|_{2}^{2}+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\langle x_{0}-\widetilde{x}_{0},z\rangle+(1-\sigma_{t}^{2})\|x_{0}^{\prime}-\widetilde{x}_{0}\|_{2}^{2}$
$\overset{\mathrm{(a)}}{\leq}-(1-\sigma_{t}^{2})\left(\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}-2\varepsilon_{0}\right)^{2}+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\langle x_{0}-\widetilde{x}_{0},z\rangle+4(1-\sigma_{t}^{2})\varepsilon_{0}^{2}$
$\overset{\mathrm{(b)}}{\leq}-(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}^{2}+4(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}\varepsilon_{0}$
$\qquad\qquad\qquad\qquad\qquad\qquad+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\left(\big{\langle}x_{j}^{*}-x_{i(x)}^{*},z\big{\rangle}+2\left(\sqrt{d}+\sqrt{C_{6}k\log k}\right)\varepsilon_{0}\right),$ (53)
where inequality (a) holds for the property of the $\varepsilon_{0}$-net, and inequality (b) comes from the definition of $z\in\mathcal{G}$. Again, recall the definition of $\mathcal{G}$ to obtain
@@
For a sufficiently large $C_{0}$, it can be easily checked that
@@
and we have
$4\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\left(\sqrt{d}+\sqrt{C_{6}k\log k}\right)\varepsilon_{0}\leq 5\sigma_{t}^{2}k\log k.$ (56)
Substituting inequalities (54), (55) and (56) into (53), and rearranging terms, we arrive at
$\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}\widetilde{x}_{0}\Big{\|}^{2}_{2}-\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}x_{0}\Big{\|}^{2}_{2}$ $\leq-\frac{1-\sigma_{t}^{2}}{2}\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}^{2}+5\sigma_{t}^{2}k\log k$
$\overset{\mathrm{(a)}}{\leq}-\frac{1-\sigma_{t}^{2}}{8}\big{\|}x_{i(x)}^{*}-x_{0}\big{\|}_{2}^{2}+5\sigma_{t}^{2}k\log k$
$\leq-\frac{C}{8}\sigma_{t}^{2}k\log k+5\sigma_{t}^{2}k\log k$
$\overset{\mathrm{(b)}}{\leq}-\frac{C}{9}\sigma_{t}^{2}k\log k,$
where (a) holds by $\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}\geq 1/2\cdot\big{\|}x_{i(x)}^{*}-x_{0}\big{\|}_{2}$, a simple property of the covering, and (b) holds provided that $C\geq C_{7}$ is large enough. Taking the above bound collectively with inequality (52) yields

### v2

@@ C.2 Proof of Lemma 11
$\overset{\mathrm{(b)}}{\leq}\frac{1}{2}\exp\left(-\frac{C_{6}}{4}k\log k\right).$ (50)
Here, we apply the union bound in inequality (a), and rely on the fact that $C_{6}$ is large enough (e.g., $C_{6}>8C_{\mathsf{cover}}^{\prime}$) in equality (b).
Regarding the set $\mathcal{G}$, we analyze it in a similar manner. First, recall some standard concentration bounds below (see, e.g., Wainwright, (2019, Chapter 2)).
@@
$\mathbb{P}(|Z_{1}|\geq t)\leq e^{-t^{2}/2}\qquad\text{and}\qquad\mathbb{P}(\|Z\|_{2}^{2}\geq\sqrt{d}+t)\leq e^{-t^{2}/2},$
@@
$\mathbb{P}(Z\notin\mathcal{G})$ $\overset{\mathrm{(a)}}{\leq}\mathbb{P}\left(\|Z\|_{2}>\sqrt{d}+\sqrt{C_{6}k\log k}\right)+\sum_{i=1}^{s}\sum_{j=1}^{s}\mathbb{P}\left(\langle x_{i}^{*}-x_{j}^{*},Z\rangle>\sqrt{C_{6}k\log k}\,\|x_{i}^{*}-x_{j}^{*}\|_{2}\right)$
@@
$\overset{\mathrm{(b)}}{\leq}\frac{1}{2}\exp\left(-\frac{C_{6}}{4}k\log k\right),$ (51)
@@
Recalling that $X_{t}\overset{\mathrm{d}}{=}\sqrt{1-\sigma_{t}^{2}}X_{0}+\sigma_{t}Z$, one can combine inequalities (50) and (51) to yield
@@
C.3 Proof of Lemma 12
For ease of presentation, we find it convenient to define the following set
$\mathcal{E}_{t}(x)\coloneqq\left\{x_{0}\,\Big{|}\,\big{\|}x_{0}-x_{i(x)}^{*}\big{\|}_{2}\geq\frac{\sigma_{t}}{\sqrt{1-\sigma_{t}^{2}}}\cdot\sqrt{Ck\log k}\right\},$ (52)
where $C_{7}>0$ is some sufficiently large constant (to be specified momentarily) and $C>C_{7}$. It thus boils down to bounding $\mathbb{P}(\mathcal{E}_{t}(x)\mid X_{t}=x)$. Denote by $q_{X_{t}|X_{0}}(\cdot|x_{0})$ the conditional density function of $X_{t}$ given $X_{0}=x$, and $q_{0}$ the measure induced by $X_{0}$. Making use of the Bayes formula, we obtain
@@ C.3 Proof of Lemma 12
where the last inequality holds by exploiting $i(x)\in\mathcal{I}$ and the definition of $\mathcal{I}$ in (46a). It then suffices to upper bound the exponential term above.
Given that $x\in\mathcal{T}_{t}$, we can decompose it as $x=\sqrt{1-\sigma_{t}^{2}}x_{0}^{\prime}+\sigma_{t}z$ for some $x_{0}^{\prime}\in\mathcal{B}_{i(x)}$ and $z\in\mathcal{G}$. For any $x_{0}\in\mathcal{E}_{t}(x)$ and $\widetilde{x}_{0}\in\mathcal{B}_{i(x)}$, suppose that $x_{0}\in\mathcal{B}_{j}$, then direct calculation gives
@@
$\qquad=-(1-\sigma_{t}^{2})\|x_{0}^{\prime}-x_{0}\|_{2}^{2}+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\,\langle x_{0}-\widetilde{x}_{0},z\rangle+(1-\sigma_{t}^{2})\|x_{0}^{\prime}-\widetilde{x}_{0}\|_{2}^{2}$
$\qquad\overset{\mathrm{(a)}}{\leq}-(1-\sigma_{t}^{2})\left(\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}-2\varepsilon_{0}\right)^{2}+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\langle x_{0}-\widetilde{x}_{0},z\rangle+4(1-\sigma_{t}^{2})\varepsilon_{0}^{2}$
$\qquad=-(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}^{2}+4(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}\varepsilon_{0}$
$\qquad\qquad\qquad\qquad\qquad+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\left(\big{\langle}x_{j}^{*}-x_{i(x)}^{*},z\big{\rangle}+\big{\langle}x_{0}-x_{j}^{*},z\big{\rangle}-\big{\langle}\widetilde{x}_{0}-x_{i(x)}^{*},z\big{\rangle}\right)$
$\qquad\overset{\mathrm{(b)}}{\leq}-(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}^{2}+4(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}\varepsilon_{0}$
$\qquad\qquad\qquad\qquad\qquad+2\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\left(\big{\langle}x_{j}^{*}-x_{i(x)}^{*},z\big{\rangle}+2\left(\sqrt{d}+\sqrt{C_{6}k\log k}\right)\varepsilon_{0}\right),$ (54)
where inequality (a) holds due to the property of the $\varepsilon_{0}$-net and the triangle inequality, and inequality (b) comes from the definition of $z\in\mathcal{G}$ as well as Cauchy-Schwarz. To further bound (54), recalling the definition of $\mathcal{G}$ again gives
@@
In addition, for a sufficiently large constant $C_{0}>0$, it is readily seen that
@@
where we recall that $d\lesssim\mathrm{poly}(k)$, $\delta=\Omega(1/\mathrm{poly}(k))$ and $\sigma_{\delta}=\sqrt{1-e^{-2\delta}}\asymp\min\{\delta,1\}$. Moreover,
$4\sqrt{\sigma_{t}^{2}(1-\sigma_{t}^{2})}\left(\sqrt{d}+\sqrt{C_{6}k\log k}\right)\varepsilon_{0}\leq 5C_{6}\sigma_{t}^{2}k\log k$ (57)
for large enough $C_{6}>0$. Substituting (55), (56) and (57) into (54), and rearranging terms, we arrive at
$\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}\widetilde{x}_{0}\Big{\|}^{2}_{2}-\Big{\|}x-\sqrt{1-\sigma_{t}^{2}}x_{0}\Big{\|}_{2}^{2}$
$\leq-(1-\sigma_{t}^{2})\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}^{2}+\left(4(1-\sigma_{t}^{2})\varepsilon_{0}+2\sqrt{C_{6}\sigma_{t}^{2}(1-\sigma_{t}^{2})k\log k}\right)\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}+5C_{6}\sigma_{t}^{2}k\log k$
$\overset{\mathrm{(a)}}{\leq}-\frac{1-\sigma_{t}^{2}}{2}\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}^{2}+5C_{6}\sigma_{t}^{2}k\log k$
$\overset{\mathrm{(b)}}{\leq}-\frac{1-\sigma_{t}^{2}}{8}\big{\|}x_{i(x)}^{*}-x_{0}\big{\|}_{2}^{2}+5C_{6}\sigma_{t}^{2}k\log k$
$\overset{\mathrm{(c)}}{\leq}-\frac{C}{8}\sigma_{t}^{2}k\log k+5C_{6}\sigma_{t}^{2}k\log k$
$\overset{\mathrm{(d)}}{\leq}-\frac{C}{9}\sigma_{t}^{2}k\log k,$
where (a) is from the definition (52) and the fact that $C_{0}$ and $C$ are large enough, (b) results from $\big{\|}x_{i(x)}^{*}-x_{j}^{*}\big{\|}_{2}\geq\big{\|}x_{i(x)}^{*}-x_{0}\big{\|}_{2}/2$, a basic property of the construction of the covering, (c) makes use of the definition of $\mathcal{E}_{t}$ in (52), and (d) holds provided that $C$ is large enough. Taking the above bound collectively with inequality (53) yields